# Segmentation and Smoothing Affect Explanation Quality More Than the Choice of Perturbation-based XAI Method for Image Explanations

### v2

@@ III Results and Analysis
The results consist of the LIF, MIF, and SRG metrics for every attribution pipeline. As this is too much data to present in this work, it is summarized as the average SRG metric for different parameter combinations. The tables show for some parameter combinations the SRG in % averaged over all pipelines with those combinations (i.e., the average result for all combinations of the unspecified parameters). The complete data is available online [^3] https://bit.ly/smooth-mask-perturbation.
The average SRG for pipelines with different sampling methods and sample sizes over the different attribution methods is shown in Table II. Unsurprisingly, increasing sample size yields improved performance. What is surprising is that random sampling significantly outperforms entropic and does so even for SHAP, for which entropic is the default sampling method. It is noteworthy that the attribution method (row-wise) in general has less impact on the SRG value than the sampling method (column-wise). Furthermore, some attribution methods seem to fail when used together with certain sampling methods and sample sizes. For non-random sampling with a sample size of 50 all attribution methods have the same performance. Under these circumstances, CIU, PDA, and RISE mathematically produce the same rankings of segments to which LIME and SHAP consistently converge.
@@
The results of different combinations of segmenting, perturbing, and attribution can be found in Table III. Again, the attribution calculation method has less impact on performance compared to the other parameters. Notably, average performance increases when attributing per-pixel for all combinations of segmenting, perturbing, and attribution methods. The Gaussian filter does not give a significantly different performance for grid segmentation compared to bilinear upsampling. However, as the Gaussian filter was implemented to mimic bilinear upsampling in this case, the results show that there is no harm in choosing one over the other. Instead, the benefit of smoothing with a Gaussian filter becomes obvious in the significant performance increase seen with SLIC over grid segmentation. As SLIC with pixel attribution is not possible with bilinear upsampling, this suggests that Gaussian filters enable the use of better segmentation methods without sacrificing performance on grid segmentation.
@@
Segmenting + Perturbing Attribute per CIU PDA LIME SHAP RISE
@@ IV Discussion
Most works that introduce or compare perturbation-based image explanation methods do not examine the parameters of the explanation pipelines separately. This leads to a poor understanding of why a method performs well and can lead to questionable claims of some methods being better than others [28, 29]. The results of this work show that each parameter of the perturbation-based image explanations pipeline significantly impacts performance.
From the results, it is clear that using SLIC as opposed to grid segmenting improves performance significantly. Per-pixel attribution, which requires smooth masks, also improves performance compared to per-segment attribution. Finally, using SLIC with per-pixel attribution further improved performance. This shows that using Gaussian filters or other alternatives to bilinear smoothing is promising as it allows the use of better-performing segmentation methods. The choice of attribution calculation has a noticeably smaller impact on performance when compared to the choice of segmentation method or the difference between per-segment and per-pixel attribution. It is interesting that PDA and RISE attribution perform differently when they perform equally in the ideal case. It seems that, with a limited sample size, using the average decrease in prediction when a segment is excluded (PDA) is less stable than using the average increase when it is (RISE).
@@
It is clear that an increased sample size does improve performance. However, the computation needed scales directly with the sample size. With ever-increasing computational demands by newer DNN models, even a low sample size let alone thousands of samples, may be unrealistic to presume for an explanation of a single decision. Developing perturbation-based methods that can give good explanations with low sample size is therefore a promising future direction. In some cases, such as medical diagnosis prediction, the need for and the value of explanations are likely high enough that it is worth increasing computational demands by factors of thousands.
@@
The evaluation in this work relied on the explanation methods being separable into different parameters that could be combined in various ways. This is not always the case, even if the method otherwise produces sound explanations. For example, the original RISE implementation shifts the perturbation masks by some pixels so as not to center the same pixels every time. This approach works with the RISE attribution calculation since it can directly assign influence to pixels. However, this is not feasible for other methods, as such shifting has not been used in this evaluation. Additionally, the use of SRG and similar metrics requires attribution scores for individual features. For example, the CIU method can use intermediate concepts to explain combinations of features, which results in explanations for the combinations rather than the individual features.
@@
This work shows how different parameters affect performance as measured by SRG. SRG measures how well the explanation reflects the actual impact of the attributed regions on the model. While ensuring that attribution maps explain the model correctly is helpful, it is not known whether it correlates with usefulness to humans. Existing evaluation on whether proxy metrics correspond to usefulness for humans is limited and requires further investigation [13]. It is also questionable whether proxy metrics can be used to compare different types of attribution maps at all. Attribution maps that indicate large regions, that provide heatmaps, or that show important edges and textures are likely to be interpreted differently by humans in a way not reflected by most proxy metrics.

### v3

@@ III Results and Analysis
The results consist of the LIF, MIF, and SRG metrics for every attribution pipeline. As this is too much data to present in this work, it is summarized as the average SRG metric for different parameter combinations. The tables show for some parameter combinations the SRG in % averaged over all pipelines with those combinations (i.e., the average result for all combinations of the unspecified parameters).
The average SRG for pipelines with different sampling methods and sample sizes over the different attribution methods is shown in Table II. Notably, sampling has a larger impact on performance than the attribution calculation. Unsurprisingly, increasing sample size yields improved performance. What is surprising is that random sampling significantly outperforms entropic sampling, even for SHAP, for which entropic sampling is the default sampling method. This may be caused by the manner in which entropic sampling was implemented in this work as described later. Furthermore, some attribution methods seem to fail when used together with certain sampling methods and sample sizes. LIME and SHAP perform poorly when using 50 random samples, likely due to having too little feature-specific information for least squares to have a single useful convergence. PDA also performs poorly when using entropic sampling at higher sample sizes, the reason for which will be described later in this section. For non-random sampling with a sample size of 50, all attribution methods have the same performance. Under these circumstances, CIU, PDA, and RISE mathematically produce the same rankings of segments to which LIME and SHAP consistently converge.
@@
The results of different combinations of segmenting, perturbing, and attribution can be found in Table III. Again, the attribution calculation has less impact on performance compared to the other parameters. Notably, average performance increases when attributing per-pixel for all combinations of segmenting, perturbing, and attribution methods. The Gaussian filter does not give a significantly different performance for grid segmentation compared to bilinear upsampling. However, as the Gaussian filter was implemented to mimic bilinear upsampling in this case, the results show that there is no harm in choosing one over the other. Instead, the benefit of smoothing with a Gaussian filter becomes obvious in the significant performance increase seen with SLIC over grid segmentation. As SLIC with pixel attribution is not possible with bilinear upsampling, this suggests that Gaussian filters enable the use of better segmentation methods without sacrificing performance on grid segmentation.
@@
Segmenting +
Perturbing
Attribute
per
CIU PDA LIME SHAP RISE
@@
It is also informative to inspect attribution maps generated for some example images that have been used in the evaluation. The attribution maps generated are overlaid on the original image to make it easy to see which part of the image has a higher attribution than another. The pixels are colored in an even spectrum between red and blue according to their rank from most attribution to least. This is in contrast to how it is commonly done, where each pixel is colored according to the value of its attribution relative to the least and greatest attribution in the image. The reason for this approach is that it is easier to assess the order in which pixels are occluded in the SRG metric, which does not care for the distribution of attribution, only the order.
Figure 3: Attribution maps per attribution method overlaid an image from the ostrich class explaining the correct prediction by AlexNet. The maps are attributed per-pixel and use grid+bilinear segmentiation and entropic sampling with a sample size of 400. The pixels are colored in an even spectrum between red and blue according to their rank from most attribution to least.
Fig. 3 compares the attribution maps generated using the PDA, LIME, SHAP, and RISE attribution methods. The comparison uses entropic sampling with a sample size of 400 to investigate a scenario where the performance between the methods varies. Aside from PDA, it is clear that the methods generate very similar attribution maps. This behavior is consistent between images and sample sizes. By switching to random sampling, PDA will generate attribution maps that are similar to those of the other methods. This behavior of PDA, seemingly arbitrarily giving less attribution to the pixels at the top of the image can be used to explain its poor performance with entropic sampling. The way entropic sampling is performed in this work is the cause of this behavior, as it alternates between all combinations where $x$ features are perturbed and all combinations where all except $x$ features are perturbed, with $x$ increasing from 0. Since 400, 4000, and 8000 are not neat cutoff points, this systematically leaves some segments being perturbed more often than others. This affects both PDA and RISE as they respectively use the average output when the feature is and isn’t perturbed. However, since the sampling starts with fewer samples being perturbed, this results in the difference in the number of times each feature is perturbed being higher (ranging from 56 to 98 for sample size 400) than the difference in the number of times it isn’t perturbed (ranging from 302 to 344). This means that PDA is much more severely affected by the imbalance in sampling than RISE is. To verify this hypothesis, the experiment has been conducted again with the reversed alternation of perturbing and not perturbing features in entropic sampling. The results of these additional experiments can be found in Table IV. When alternation is reversed, RISE instead suffers from a similar loss of performance as PDA does, while PDA gains a significant boost to performance. This seems to verify the hypothesis that PDA and RISE are sensitive to unbalanced sampling.
TABLE IV: The SRG in % for PDA and RISE pipelines with entropic sampling where the alternation between perturbing $x$ samples and perturbing all but $x$ samples has been reversed. The results are averaged for all pipelines with the specified parameters.
Attribution
Sampling Sample size PDA RISE
Entropic (reverse alternation) 400 19.3 9.5
Entropic (reverse alternation) 4000/8000 20.5 16.4
Figure 4: Per-segment and per-pixel attribution maps generated using both Grid+Gaussian and SLIC+Gaussian segmentation overlaid an image of a snail, explaining the correct prediction by AlexNet. The maps were generated using "only one" sampling and are equivalent across all attribution methods. The pixels are colored in an even spectrum between red and blue according to their rank from highest attribution to lowest.
Fig. 4 compares the per-segment and per-pixel attribution maps generated with grid and SLIC segmentation. The comparison uses Gaussian smoothing for both segmentation methods for a fairer comparison and "only one" sampling to use a case where all attribution methods give the same attribution maps. The figure highlights one advantage of per-pixel attribution: its ability to smooth the borders between segments means that shapes that do not align with the segments can be better contained in the areas of high or low attribution without a harsh cutoff. This advantage is more prevalent when using grid segmentation, as the method does not take the shapes of the image into account when segmenting. SLIC has a similar benefit over grid segmentation as its segments are adapted to the actual features of the image. That SLIC already fits the features of the image is likely why it gets relatively less improvement from per-pixel attribution compared to grid segmentation. In the case of this image, the difference in the per-pixel case is essentially negligible.
@@ IV Discussion
Most works that introduce or compare perturbation-based image explanation methods do not examine the parameters of the explanation pipelines separately. This leads to a poor understanding of why a method performs well and can lead to questionable claims of some methods being better than others [velmurugan2023through, miro2024assessing]. The results of this work show that each parameter of the perturbation-based image explanations pipeline significantly impacts performance.
From the results, it is clear that using SLIC as opposed to grid segmenting improves performance significantly. Per-pixel attribution, which requires smooth masks, also improves performance compared to per-segment attribution. Finally, using SLIC with per-pixel attribution further improved performance. This shows that using Gaussian filters or other alternatives to bilinear smoothing is promising, as it allows the use of better-performing segmentation methods. The choice of attribution calculation has a noticeably smaller impact on performance when compared to the choice of segmentation method or the difference between per-segment and per-pixel attribution. The largest impact on performance came from issues that PDA and RISE attribution encounter when sampling is biased towards certain features. It is likely that using more careful sampling methods would eliminate these differences, making the choice of attribution method seemingly arbitrary when it comes to occlusion metrics. This is supported by visualizations of the attribution maps generated for the images, which are similar. When it comes to feature attribution for images, it is arguable that performance is not the most relevant criterion for choosing an attribution method. Instead, it is useful to consider other factors when choosing an attribution method. CIU, PDA, and RISE are all likely easier to explain to a lay user than LIME and SHAP, which is an advantage. CIU’s ability to differentiate between importance and influence is specific to that attribution method. On the other hand, CIU, in its current form, cannot assign attribution to single features when using sampling methods other than "only one" and "all but one", which limits its performance on occlusion metrics in high sampling situations.
@@
It is clear that an increased sample size does improve performance. However, as model calls dominate the computational demands of the other parts of the pipelines, the total computation of each pipeline scales directly with sample size. With ever-increasing computational demands by newer DNN models, even a low sample size, let alone thousands of samples, may be unrealistic to presume for an explanation of a single decision. In some cases, such as medical diagnosis prediction, the need for and the value of explanations are likely high enough that it is worth increasing computational demands by factors of thousands, but for many other cases, low-computation explanations are needed. Developing perturbation-based methods that can give good explanations with a low sample size is, therefore, a promising future direction. One approach to this could be by using a limited number of samples initially and using the results from those samples to determine the next set of samples. Such an approach would essentially make the sampling, perturbing, and predicting steps of the pipelines recurrent.
@@
The evaluation in this work relied on the explanation methods being separable into different parameters that could be combined in various ways. This is not always the case, even if the method otherwise produces sound explanations. For example, the original RISE implementation shifts the perturbation masks by some pixels so as not to center the same pixels every time. This approach works with the RISE attribution calculation since it can directly assign influence to pixels. However, this is not feasible for other methods, as such shifting has not been used in this evaluation. Additionally, the use of SRG and similar metrics requires attribution scores for individual features. For example, the CIU method can use intermediate concepts to explain combinations of features, which results in explanations for the combinations rather than for the individual features.
@@
This work shows how different parameters affect performance as measured by SRG. SRG measures how well the explanation reflects the actual impact of the attributed regions on the model. While ensuring that attribution maps explain the model correctly is helpful, it is not known whether it correlates with usefulness to humans. Existing evaluations on whether proxy metrics correspond to usefulness for humans are limited and require further investigation [nguyen2021effectiveness]. It is also questionable whether proxy metrics can be used to compare different types of attribution maps at all. Attribution maps that indicate large regions, that provide heatmaps, or that show important edges and textures are likely to be interpreted differently by humans in a way not reflected by most proxy metrics.
Finally, all the steps of the perturbation-based image attribution pipeline need to be considered when implementing an explanation system. Future research should be careful when comparing different pipelines with each other, not to assign performance increases to some individual part of a pipeline unless the other parameters are the same. Similar pipelines likely exist in other domains beyond images, which would likely also benefit from further investigation.